### v1

# A note on Johnson’s rule for minimizing makespan in the Two-Machine Flow Shop scheduling problem

Abstract
We consider Johnson’s rule for minimizing the makespan in the two-machine flow shop scheduling problem. We show that although the worst-case complexity of Johnson’s rule is $O(n\log n)$, since it requires a complete sorting of the jobs, it is possible to detect in linear time whenever a full sort can be avoided and the optimal solution can be computed in linear time. Computational testing indicates that the linear-time complexity always occurs in practice on standard benchmark instances with uniform distribution of the processing times.
1 Introduction
In this short note, we consider the classical two-machine flow shop scheduling problem with the objective of minimizing the makespan, commonly denoted by $F2\;||\;C_{\max}$ [3]. A set $J=\{1,\dots,n\}$ of jobs must be processed on two machines $M_{1}$ and $M_{2}$ in this order (flow shop). Job $j$ requires processing time $p_{1,j}$ on $M_{1}$ and $p_{2,j}$ on $M_{2}$. All jobs are available at time 0, each machine can handle at most one job at a time and preemption is not allowed. It is well known that there always exists an optimal schedule which is a permutation schedule and can be obtained by the following Johnson’s rule [4] that first schedules the set $A$ of jobs such that $p_{1,j}\leq p_{2,j}$ in nondecreasing of $p_{1,j}$ and then the set $B$ of jobs such that $p_{1,j}>p_{2,j}$ in nonincreasing of $p_{2,j}$ .
Algorithm 1 Johnson’s rule for $F2\|\!C_{\max}$
1:INPUT - Processing times $\{(p_{1j},p_{2j})\}_{j=1}^{n}$
[…]
3:Sort AA by nondecreasing $p_{1j}$; sort $B$ by nonincreasing $p_{2j}$
4:return sequence $\pi\leftarrow$ Concatenate(A,B)(A,B)
Due to the required sorting, Johnson’s rule runs in $O(n\log n)$ time and $O(n)$ space. As mentioned in [7], schedules obtained by Johnson’s rule ”are by no means the only schedules that are optimal for $F2\;||\;C_{\max}$. The class of optimal schedules appears to be hard to characterize and data dependent.” Here, we make a step in this direction, by showing that a simple procedure requiring linear time allows to check whether or not a full sort can be avoided and, correspondingly, an optimal solution can be computed in linear time. In that case, a huge set of equivalent optimal sequences is also detected.
The rest of this note proceeds as follows. In Section 2 we show how to detect in linear time subsets of jobs that do not need to be sorted in the search of an optimal solution as any permutation of those jobs would also induce an optimal sequence. Correspondingly, it is possible to verify whether the size of the remaining jobs to be sorted is small enough to guarantee an overall linear time complexity. As a counterpart, we also present a family of instances where a full sort is necessary. Then, in Section 3, computational testing indicates that, on standard uniform distributions, for instances with size $n=100,1000,10000,100000$ and $1000000$, an optimal solution was always reached in linear time. Finally, Section 4 concludes the paper with final remarks.
2 Revisiting Johnson’s rule for problem $F2\;||\;C_{\max}$
Let assume, without loss of generality, that there are $n=n_{A}+n_{B}$ jobs that are indexed according to Johnson’s sequence where jobs $1,\dots,n_{A}$ belong to set $A$ and jobs $n_{A}+1,\dots,n_{B}$ belong to set $B$. Let also denote by $p^{\max}_{A}$ ($p^{\max}_{B}$) the largest processing time in subset AA ($B$). Then, let $P_{1}=\sum_{i=1}^{n}p_{1,i}$ be the total processing time on machine $M_{1}$. Similarly, let $P_{2}=\sum_{i=1}^{n}p_{2,i}$ be the total processing time on machine $M_{2}$. Correspondingly, let $P_{1A}=\sum_{i=1}^{n_{A}}p_{1,i}$ ($P_{2A}=\sum_{i=1}^{n_{A}}p_{2,i}$) be the total processing time of the jobs in subset $A$ on machine $M_{1}$ ($M_{2}$) and let $P_{1B}=\sum_{i=n_{A}+1}^{n_{B}}p_{1,i}$ ($P_{2B}=\sum_{i=n_{A}+1}^{n_{B}}p_{2,i}$) be the total processing time of the jobs in subset $B$ on machine $M_{1}$ ($M_{2}$). Finally, when jobs are processed on the machines, for each job $i$, let denote by $C_{1,i}$ ($C_{2,i}$) the related completion time on machine $M_{1}$ ($M_{2}$).
We remark that the $F2\;||\;C_{\max}$ problem possesses the reversibility property presented below. Consider two $2$-machine flow shops where $p^{(1)}_{i,j}$ and $p^{(2)}_{i,j}$ denote the processing times of job $j$ on machine $i$ in the first and second flow shop, respectively.
Property 1.
[6, 7] If $p^{(1)}_{1,j}=p^{(2)}_{2,j}$ and $p^{(1)}_{2,j}=p^{(2)}_{1,j}$, then sequence $j_{1},\dots,j_{n}$ in the first flow shop results in the same makespan as sequence $j_{n},\dots,j_{1}$ in the second flow shop.
The following properties provide a sufficient condition for finding an optimal solution of problem $F2\;||\;C_{\max}$ without requiring a full sort of the job set.
Property 2.
If $P_{1A}\leq P_{2A}-p^{\max}_{A}$, let $k_{A}$ denote the smallest index such that
$\sum_{i=1}^{k_{A}}p_{1,i}\leq\sum_{i=1}^{k_{A}}p_{2,i}-p^{\max}_{A}.$
Then, for set $A$, an optimal schedule is composed by the subsequence of jobs $1,\dots,k_{A}$, followed by jobs $k_{A}+1,\dots,n_{A}$ sequenced in any order.
Proof.
As $p_{1,i}\leq p_{2,i}\;\forall i\in A$, once
$\sum_{j=1}^{k_{A}}p_{1,j}\leq\sum_{i=1}^{k_{A}}p_{2,j}-p^{\max}_{A},$
for each $i\in\{k+1,\dots,n_{A}\}$, whatever job $i^{*}$ is selected for position $k+1$, the completion time on the first machine of that job respects the inequality
$C_{1,i^{*}}=\sum_{j=1}^{k_{A}}p_{1,j}+p_{1,i^{*}}\leq\sum_{j=1}^{k_{A}}p_{1,j}+p^{\max}_{A}\leq\sum_{j=1}^{k_{A}}p_{2,j}$
namely no idle time is enforced on machine $M_{2}$ for job $i^{*}$. Furthermore, the gap between the completion times on the two machines for the job in positions $\geq k+1$ can only increase with respect to the gap of the job in position $k$, and correspondingly no idle time can ever occur on machine $M_{2}$ for positions $k+1,\dots,n_{A}$. Consequently, the completion time of the last job of set $A$ on the second machine will be the same regardless of the permutation considered for the jobs in positions $k+1,\dots,n_{A}$. ∎
Property 3.
If $P_{2B}\leq P_{1B}-p^{\max}_{B}$, let $k_{B}$ denote the largest index such that
$\sum_{i=k_{B}}^{n_{B}}p_{2,i}\leq\sum_{i=k_{B}}^{N_{B}}p_{1,i}-p^{\max}_{B}.$
Then, for set $B$, an optimal schedule is composed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order followed by subsequence $k_{B},\dots,n_{B}$ .
Proof.
Due to the reversibility property 1, the conditions of Property 3 correspond to the conditions of Property 2 applied to the reversible flow shop problem. ∎
Combining properties 2, 3, the following property allows to detect in linear time whether a sufficient condition for solving problem $F2\;||\;C_{\max}$ in linear time also holds.
Property 4.
If Properties 2, 3 hold with $k_{A}\log k_{A}\leq n$ and $(n-k_{B}+1)\log(n-k_{B}+1)\leq n$, then the optimal solution of problem $F2\;||\;C_{\max}$ can be computed in linear time where only jobs $1,\dots k_{A}$ need to be sorted in subset $A$ and only jobs $k_{B},\dots,n$ need to be sorted in subset $B$.
Proof.
Splitting the job set into subsets $A$ and $B$, clearly takes linear time. Finding job $k_{A}$ in set $A$ and job $k_{B}$ in set $B$ can be done in linear time according to the standard technique of median finding [1] (in a similar way it is used in the search of the critical item for finding the optimal continuous relaxation for the $0/1$ knapsack problem [2, 5]). But then, if $k_{A}\log k_{A}\leq n$ and $(n-k_{B}+1)\log(n-k_{B}+1)\leq n$ hold, sorting job sets $1,\dots k_{A}$ and $k_{B},\dots,n$ also requires linear time. ∎
Job $j$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
$p_{1,j}$ 1 2 3 4 5 6 7 8 9 10 9 8 8 8 7 7 9 10
$p_{2,j}$ 8 9 7 8 9 7 9 10 10 9 7 7 6 5 4 3 2 1
[…]
Figure 1: 18-job instance with $\sum p_{1,i}=\sum p_{2,i}=121$: Gantt diagram of Johnson’s schedule and of another equivalent optimal schedule
Consider the instance in Figure $1$ with $18$ jobs where Properties 2 and 3 hold with $k_{A}=n-k_{B}=2$. Correspondingly, in the optimal solution, for subset $A$ only the first two jobs, colored in gray, need to be sorted and placed at the beginning of the sequence. Similarly, for subset $B$ only the last two jobs, also colored in gray, need to be sorted and placed at the end of the sequence.
For the remaining jobs $\{3,4,5,6,7,8,9\}$ in subset $A$, colored in green, any permutation yields an optimal solution. The same consideration holds for the remaining jobs $\{10,11,12,13,14,15,16\}$ in subset $B$ also colored in blue. Figure $1$ shows also the corresponding Gantt chart for Johnson’s sequence and for an equivalent optimal sequence in which the order of jobs $\{3,4,5,6,7,8,9\}$ and $\{10,11,12,13,14,15,16\}$ is reversed.
An even stronger result in terms of equivalent optimal schedules, can be deduced by the following properties 5, 6 whenever either $P_{1}\leq P_{2}-p^{\max}$, or $P_{2}\leq P_{1}-p^{\max}$ holds.
Property 5.
If $P_{1}\leq P_{2}-p^{\max}$, let $k_{A}$ denote the smallest index such that
$\sum_{i=1}^{k_{A}}p_{1,i}\leq\sum_{i=1}^{k_{A}}p_{2,i}-p^{\max}_{A}.$
Then, an optimal schedule is composed by sequencing initially set $A$ with first the subsequence of jobs $1,\dots,k_{A}$, followed by jobs $k_{A}+1,\dots,n_{A}$ sequenced in any order followed by set $B$ sequenced in any order. Also, if $k_{A}\log k_{A}\leq n$, then the optimal solution of problem $F2\;||\;C_{\max}$ can be computed in linear time.
Proof.
The first part of the proof related to set $A$ trivially holds from Property 2 as $p^{\max}_{A}\leq p^{\max}$. For the second part, as
$P_{1}=P_{1A}+P_{1B}\leq P_{2}-p^{\max}=P_{2A}+P_{2B}-p^{\max},$
we have
$P_{2A}-P_{1A}\geq P_{1B}-P_{2B}+p^{\max}.$
Correspondingly, whatever schedule on set $B$, the makespan will be always given by $C_{2,N_{A}}+P_{2B}$, that is no idle time will be induced on machine $2$ in set $B$. For the time complexity, the same analysis of Property 4 holds. ∎
Property 6.
If $P_{2}\leq P_{1}-p^{\max}$, let $k_{B}$ denote the argest index such that
$\sum_{i=k_{B}}^{n_{B}}p_{2,i}\leq\sum_{i=k_{B}}^{N_{B}}p_{1,i}-p^{\max}_{B}.$
Then, an optimal schedule is composed by scheduling initially set $A$ sequenced in any order, followed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order, followed by subsequence $k_{B},\dots,n_{B}$ . Also, if $(n-k_{B}+1)\log(n-k_{B}+1)\leq n$, then the optimal solution of problem $F2\;||\;C_{\max}$ can be computed in linear time.
Proof.
Due to the reversibility property 1, the conditions of Property 6 correspond to the conditions of Property 5 applied to the reversible flow shop problem. ∎
As a counterpart, we show that, in the worst case, a full sort of both subsets $A$ and $B$ may occur. Indeed, consider a $2$-machine flow shop problem with $2n$ jobs. Here, the first $n$ jobs belong to subset $A$ and the processing times are as follows: $\forall j\in 1,\dots,n$, $p_{1,j}=j<p_{2,j}=j+1$. Then, the second $n$ jobs belong to subset $B$ and the processing times are as follows: $\forall j\in n+1,\dots,2n$, $p_{1,j}=2n-j+2>p_{2,j}=2n-j+1$. It is straightforward to show that the unique optimum is given by sequence $1,2,\dots,n-1,n,n+1,\dots 2n-1,2n$ and thus a complete sorting of subset $A$ (in nondecreasing processing time order on the first machine) and subset $B$ (in nonincreasing processing time order on the second first machine) is necessary. In figure 2, the related instance and relevant unique optimal solution for $n=4$ with makespan $25$ is presented.
Job $j$ 1 2 3 4 5 6 7 8
$p_{1,j}$ 1 2 3 4 5 4 3 2
$p_{2,j}$ 2 3 4 5 4 3 2 1
[…]
Figure 2: An instance with unique optimal sequence
3 Computational testing
In flow shop standard instances, the processing times are typically generated according to uniform distributions , see e.g. [8]. Correspondingly, in order to check what is in practice the size of $k_{A}$ and $k_{B}$, we considered instances with sizes $10^{2},10^{3},10^{4}$ and $10^{5}$ and processing times were generated according to uniform distributions $1\dots 10^{2}],[1\dots 10^{3}],[1\dots 10^{4}]$ and $1\dots 10^{5}]$. For each pair (size, distribution), $10$ instances were generated, for a total of 160 instances. The results are summarized in Table $1$. The first column provides the number of jobs. The second column indicates the maximum processing time of the related uniform distribution. The third (fourth) column indicates the maximum value of $k_{A}$ ($k_{B}$) reached among the $10$ generated instances of the same pair (size, distribution). The fifth column indicates for each pair pair (size, distribution) the number of times Property 5 or Property 6 hold. The last column indicates the number of times an optimal sequence can be computed in $O(n)$ time.
From the Table, we evince that, in all instances, both $k_{A}$ and $k_{B}$ where not superior to $6$, meaning that both sets $A$ and $B$ required a limited sorting involving up to $6$ jobs. Correspondingly, for all these instances, the optimal solution of the related $F2\;||\;C_{\max}$ problem could be computed in linear time. Also, for most of the instances ($144/160$) either Property 5 or Property 6 happened to hold, that is just $k_{A}$ or $k_{B}$ jobs needed to be sorted.
njob $p_{max}$ maxmax $k_{A}$ maxmax $k_{B}$ #\# Prop 5,6 #\# $O(n)$
100 100 4 3 5 10
100 1000 3 4 8 10
100 10000 5 4 9 10
100 100000 3 3 7 10
1000 100 5 5 10 10
1000 1000 4 4 10 10
1000 10000 3 4 9 10
1000 100000 3 5 9 10
10000 100 3 6 10 10
10000 1000 4 4 10 10
10000 10000 6 4 8 10
10000 100000 4 6 10 10
100000 100 3 5 10 10
100000 1000 4 4 10 10
100000 10000 5 4 10 10
100000 100000 4 4 9 10
Table 1: Testing instances with uniform distribution
4 Conclusions
We revisited Johnson’s rule for makespan minimization in the two-machine flow shop problem. We provided a linear time sufficient condition to detect whether the optimum is computable in linear time and showed that, even if instances exist where a complete sorting of subsets $A$ and $B$ is necessary, in practice, on standard uniform distribution, only a very partial sorting is required and therefore an optimal sequence could always be computed in linear time. We remark, also, that all permutations of the jobs in $A$ and $B$ that need not to be sorted can be part of an optimal schedule inducing correspondingly huge subsets of equivalently optimal solutions.